OpenStudy (anonymous):

what is the value of 10c6

OpenStudy (anonymous):

binomial(10, 6) = (10!)/(4! 6!)= 210

OpenStudy (anonymous):

thanks where did you get the 4

OpenStudy (anonymous):

I simplified the binomial of (10,6)

OpenStudy (anonymous):

Or this way might be more clear 10C6 = 10! / (6! * 4!) = 10 * 9 * 8 * 7 / 4! = 5040/24 = 210.

OpenStudy (anonymous):

oh ok thanks that helped alot